# Bitcoin Under Volatile Block Rewards: How Mempool Statistics Can Influence Bitcoin Mining

### v1

@@ 3.1. Non-profitability of Pre-DAM Selfish Mining Under the Fixed Block Reward Model
In Bitcoin, a difficulty epoch is defined as the period during which 2016 blocks are appended to the canonical chain. At the end of each epoch, the Difficulty Adjustment Mechanism (DAM) recalibrates the difficulty target for the subsequent epoch based on the hash power estimated from the previous epoch. Under a fixed block reward model, selfish mining cannot be more profitable than honest mining during the first epoch and before the DAM is applied (Grunspan and Pérez-Marco, 2018). To illustrate this, it is essential to compare the adversary’s time-averaged profit from both honest and selfish mining prior to the difficulty adjustment. Let $e$ denote an epoch and $\lambda$ denote the mining rate, assuming all mining nodes, including the adversary, mine honestly. In this case, $\frac{1}{\lambda}$ represents the average number of rounds required to add a new block to the canonical chain. We define the adversary’s total block generation rate as the average number of rounds in which a new adversarial block is mined. We also define the adversary’s canonical block generation rate as the average number of rounds in which a new adversarial block is added to the canonical chain. Under honest mining, both the total and canonical block generation rates are the same. In the fixed-block reward model, we assume the reward for each block in the canonical chain is constant, denoted by $R$. We ignore the mining cost for simplicity, as it is the same under both honest and selfish mining.
If the adversary $\mathcal{A}$, with a stake share of $\alpha$, mines honestly during epoch $e$, its canonical block generation rate would be equal to $\lambda\alpha$, resulting in a time-averaged profit of $\lambda\alpha R$. If the adversary $\mathcal{A}$ starts selfish mining at the beginning of epoch $e$, its canonical block generation rate would be less than or equal to that under honest mining ($\leq\lambda\alpha$). This is because the difficulty of epoch $e$ is already set before the start of the epoch, and the adversary’s strategy during the epoch cannot influence the epoch difficulty. Since both the epoch difficulty and the adversary’s mining share are fixed, the adversary’s total block generation rate remains the same under both honest and selfish mining. However, under selfish mining, some of the adversarial blocks may become orphaned, leading to a decrease in the adversary’s canonical block generation rate. Therefore, the time-averaged profit from selfish mining during epoch $e$ (before a DAM is applied) is $\leq\lambda\alpha R$. Intuitively, before a DAM is applied, the adversary cannot mine blocks faster, regardless of the strategy chosen (Sarenche et al., 2024b).
@@ 3.2. Profitability of Long-Range Selfish Mining Under the Fixed Block Reward Model
Knowing that under the fixed block reward model, selfish mining cannot be profitable during the initial epoch, the adversary needs to sustain the selfish mining attack for a longer period to achieve profitability. In Bitcoin, DAM is responsible for adjusting the block generation rate to ensure that it takes, on average, 10 minutes to mine a new block. This implies that the ideal epoch duration is two weeks. Assume the adversary starts selfish mining at the beginning of epoch $e$. As a result of selfish mining, some of the blocks (both adversarial and honest) mined during epoch $e$ may become orphaned. The generation of these orphan blocks extends the duration of epoch $e$ because to add 2016 blocks to the canonical chain, more than 2016 blocks need to be mined. At the end of epoch $e$, which exceeds two weeks, the DAM decreases the mining difficulty for the following epoch to restore the canonical block generation rate to one block per 10 minutes. This reduction in difficulty leads to a higher block generation rate for the adversary, indicating that the adversary can mine new blocks in fewer rounds (Sarenche et al., 2024b).
This implies that under the fixed block reward model, selfish mining requires at least one difficulty epoch to become profitable. In practice, the attack duration must be extended further to ensure profitability. This is because the selfish miner incurs losses during the initial epoch due to orphaned blocks, necessitating a longer attack period to compensate for these losses. Let $\texttt{BR}^{\mathcal{A}}(\pi^{\texttt{selfish}})$ denote the adversarial block ratio under selfish mining, i.e., the ratio of the number of adversarial blocks to the total number of blocks available in the canonical chain. Note that this ratio is not affected by the mining difficulty and is the same before and after applying a DAM. We normalize the two-week duration of an ideal epoch to 1 and the total reward for an epoch, $2016R$, to 1 as well. The duration of the starting epoch of the attack, namely epoch $e$, is denoted by $D_{e}$ which is greater than 1. The time-averaged profit of selfish mining in epoch $e$ can be obtained as follows:
(3) $\texttt{Profit}_{e}=\frac{\texttt{BR}^{\mathcal{A}}(\pi^{\texttt{selfish}})}{D_{e}}<\texttt{BR}^{\mathcal{A}}(\pi^{\texttt{selfish}})\enspace.$
Assume the adversary continues selfish mining in the upcoming epochs $E=\{e+1,e+2,\cdots\}$. Each epoch $e^{\prime}$ in set $E$ is preceded by an epoch that has gone under the selfish mining attack and thus has a reduced level of mining difficulty compared to epoch $e$. Since for any epoch $e^{\prime}\in E$ and its preceding epoch $e^{\prime}-1$, the effective mining power that contributes to the canonical chain is the same (both are under the selfish mining attack), the DAM at the end of epoch $e^{\prime}-1$ can properly adjust the mining difficulty for epoch $e^{\prime}$ so that the duration of epoch $e^{\prime}$ becomes 1. Therefore, the time-averaged profit of selfish mining in epoch $e^{\prime}\in E$ can be obtained as follows:
(4) $\texttt{Profit}_{e^{\prime}}=\texttt{BR}^{\mathcal{A}}(\pi^{\texttt{selfish}})\enspace.$
The equality above indicates that the time-averaged profit of selfish mining in epoch $e^{\prime}\in E$ equals the adversarial block ratio. However, since the time-averaged profit during the initial epoch of the attack is less than the adversarial block ratio, the total time-averaged profit from the start of the attack remains below the attacker’s block ratio. Over a long period of selfish mining, the time-averaged profit will eventually converge to the adversarial block ratio.
In the following, we analyze the minimum number of days required for a selfish mining attack to become profitable under the fixed block reward model, i.e., the first point at which the profit from selfish mining exceeds that of honest mining since the start of the attack. This period is referred to as the profit lag in (Grunspan and Pérez-Marco, 2023). Fig. 2 represents the minimum number of days required for a selfish mining attack to become profitable under the fixed block reward model. The figure is depicted for an adversary with different communication capabilities of $\gamma=0.5$, $0.75$, and $1$. To depict Fig. 2, we have assumed that the adversary follows the optimal selfish mining strategy introduced in (Sapirshtein et al., 2017), using which the adversary can achieve the highest possible block ratio. As can be seen in Fig. 2, the duration of the initial loss does not always decrease with an increase in the adversarial mining power, which seems counterintuitive. This is because the optimal selfish mining strategy introduced in (Sapirshtein et al., 2017) only aims to maximize the selfish miner’s block ratio in the canonical chain, indicating that this strategy optimizes the selfish miner’s profitability only after the difficulty adjustment mechanism is applied and the mining difficulty is reduced. However, as indicated in equation 3, the profitability of selfish mining in the first days of the attack depends not only on the block ratio achieved under a specific strategy but also on the duration of the initial epoch resulting from that strategy. Therefore, the optimal selfish mining strategy introduced in (Sapirshtein et al., 2017), which only considers optimizing the block ratio, is not necessarily the optimal strategy that an attacker can follow before the adjustment of the difficulty. For instance, consider a scenario where the adversarial communication capability is equal to 1, meaning the selfish miner can orphan one honest block for every adversarial block mined. In this situation, as the selfish miner’s mining power increases, more honest blocks will be orphaned during the initial epoch of selfish mining, resulting in a longer duration for this initial epoch. For detailed information regarding selfish mining profit lag, readers are referred to (Grunspan and Pérez-Marco, 2023) and (Sarenche et al., 2024b).
The main takeaway that Fig. 2 serves is that under the fixed block reward model, selfish mining takes a relatively long time (more than one ideal epoch or two weeks) to become profitable. Thus far, Bitcoin has not experienced a serious selfish mining attack, as the number of detected orphan blocks is negligible. These orphan blocks can likely be attributed to delays in publishing blocks in the network. One of the reasons that might make miners reluctant to engage in selfish mining is the relatively long initial period of the loss, which is caused by the infrequent application of the difficulty adjustment mechanism (once every two weeks). Although the infrequent DAM has some drawbacks, such as unstable block throughput during fluctuations in mining power, it can disincentivize mining power destruction attacks by prolonging the period of unprofitability.
A concerning point regarding Bitcoin’s security is that, due to the halving events, transaction fee rewards are expected to become comparable to or even exceed the protocol’s block rewards in the coming years, leading to more volatile block rewards. The volatility of the reward model is caused by the variation in transaction fees across different blocks. Under the volatile block reward model, the initial loss period of selfish mining can be shortened or even eliminated, regardless of the infrequent difficulty adjustment mechanism. Therefore, a selfish miner may see an immediate increase in its profit after starting the attack, suggesting that selfish mining could become a more significant threat to the Bitcoin network in the future.
While selfish mining under the fixed block reward model requires waiting for the difficulty adjustment to become profitable, selfish mining under the volatile block reward model can be profitable even before a DAM. Similar to the fixed block reward model, selfish mining does not increase the block generation rate before a DAM under the volatile block reward model. The key factor that makes the initial epoch of selfish mining profitable under the volatile model is that the block reward in this model is not fixed and can vary depending on the amount of fees paid by the transactions included in the block. Each block is typically filled with those transactions from the mempool that offer the highest transaction fees. The total transaction fees of a block depend on various factors [^2] Other variables such as transaction broadcasting rate, transaction fee rate, block space, etc. can also affect the total fee of a block., with block generation time playing a significant role in the profitability of selfish mining before a DAM. The block generation time is defined as the time gap between the proposal time of the last block in the canonical chain and the proposal time of the new block. There is competition among Bitcoin users seeking to have their transactions included in the blockchain quickly, requiring them to select an appropriate transaction fee. The longer it takes to generate a new block, the more intense the competition among transactions to be included in the next block becomes, resulting in a greater number of fee-valuable transactions being broadcast to the mempool. This implies that the total transaction fees included in a block increase as the block generation time grows. Let $\lambda_{1}$ [^3] Note that $\lambda_{1}=\lambda\alpha$, where $\lambda$ is the total block generation rate. and $R_{1}$ denote the adversarial canonical block mining rate and the average block reward under the volatile reward model when all the miners including the adversary mine honestly. Also, let $\lambda_{2}$ and $R_{2}$ denote the adversarial canonical block mining rate and the average block reward under the volatile reward model during the first epoch of selfish mining. As discussed in Section 3.1, the adversary’s time-averaged profit during the honest mining and the first epoch of selfish mining are equal to $\lambda_{1}R_{1}$ and $\lambda_{2}R_{2}$, respectively. By conducting a selfish mining attack, the adversarial canonical block generation rate decreases in the first epoch of the attack as some of the adversarial blocks may become orphaned, implying that $\lambda_{2}\leq\lambda_{1}$. However, unlike the fixed block reward model, $R_{2}$ is not necessarily equal to $R_{1}$. In the volatile model, as a result of the increase in the block generation time [^4] The network block generation time is inversely proportional to the total block generation rate. during the first epoch of the attack, the average block reward may increase, implying that $R_{2}\geq R_{1}$. The reason behind this increase in the block reward is intuitively illustrated in Fig. 3. Therefore, one of the time-averaged profits $\lambda_{i}R_{i}$ for $i\in\{1,2\}$ may be greater than the other depending on the adversary’s mining parameters and the mempool statistics. This shows that under the volatile reward block model, selfish mining can be profitable even in the first epoch of the attack.
3.4. Profitability of Selfish Mining After DAM Under the Volatile Block Reward Model
Let $e$ denote the initial epoch of selfish mining. As discussed in Section 3.2, after applying DAM at the end of epoch $e$, the mining difficulty for the subsequent epoch $e+1$ will be reduced. This reduction in difficulty implies that the adversary will be able to mine new blocks in a shorter time. Consequently, the average number of canonical adversarial blocks per unit of time will increase after the difficulty adjustment compared to before it. [^5] This is the case if the selfish miner has the minimum power share that makes selfish mining profitable. Regardless of whether transaction fees are comparable to the protocol reward, the increase in the average number of adversarial blocks per unit of time will raise the adversary’s reward per unit of time. Therefore, the adversary’s time-averaged profit increases after the difficulty adjustment in both the fixed and volatile reward models. An important question that arises in this context is which reward model—fixed or volatile—makes selfish mining more profitable after the difficulty adjustment. As demonstrated in the WeRLman paper (Bar-Zur et al., 2022), the minimum hash rate required for profitable deviation from honest mining after difficulty adjustment decreases in the volatile reward model. To explain how the volatile reward model lowers the deviation mining share threshold, we must focus on blocks with significantly higher rewards compared to those around them. These blocks may contain a whale transaction with an unusually large fee or may result from exceptionally long block generation times. Once such a block is mined, the adversary is incentivized to orphan the block, rather than extend it, in an attempt to steal the valuable transactions included within.
However, this reduction in the deviation mining share threshold does not necessarily imply that difficulty-adjusted selfish mining in the volatile reward setting is always more profitable for all miners and across all mempool patterns compared to the same attack under the fixed reward model. Before the difficulty adjustment, the main advantage of selfish mining in the volatile model compared to the fixed reward model is that it increases the average reward of an adversarial block after the attack is initiated. This occurs because selfish mining prolongs the canonical block generation time, allowing the next block to accumulate more valuable transaction fees. After the difficulty adjustment, however, the average block generation time in the canonical chain reverts to the standard 10-minute duration. This means that, after the difficulty adjustment, selfish mining under the volatile model loses its advantage of producing adversarial blocks with longer average generation times.
To compare difficulty-adjusted selfish mining profitability under the volatile and fixed reward models, we need to examine the ratio of the average canonical block reward before and after the attack in both models. Note that this ratio is 1 in the fixed reward model. However, in the volatile model, although the average block generation time remains the same before and after the attack, the different pattern of canonical block arrivals after the difficulty-adjusted selfish mining results in the average block reward not being necessarily the same before and after the attack. In the WeRLman paper (Bar-Zur et al., 2022), it is assumed that the rewards of all the normal blocks before and after the attack are the same; however, this assumption does not fully reflect what happens in practice. For instance, after the difficulty adjustment, the reduction in mining difficulty leads to shorter average block generation times for those canonical blocks that have no orphan blocks between them and their parent blocks, suggesting that these blocks tend to have lower fees compared to normal blocks generated before the attack.
To clarify that the average block reward under the volatile block reward model is not necessarily the same before and after difficulty-adjusted selfish mining, we first define the time-fee function and then present a theorem related to it. A time-fee function estimates the total transaction fees included in a block as a function of block generation time. Different mempool patterns can result in different time-fee functions.
Theorem 3.1.
The proof of Theorem 3.1 is presented in Appendix C.
Observing Bitcoin mempool statistics (mem, 2024) reveals that there are frequent periods during which the corresponding time-fee function of the mempool is strictly concave, meaning the rate of increase in the block fee decreases over time. In other words, in most cases, a block $B_{1}$ with twice the generation time of block $B_{2}$ will accumulate less than twice the transaction fees of $B_{2}$. According to Theorem 3.1, under a concave time-fee function, the average canonical block fees after difficulty-adjusted selfish mining are lower than those before the attack. Therefore, since the ratio of the average block reward before and after the attack under the fixed reward model is 1, the percentage increase in profitability from difficulty-adjusted selfish mining under the fixed block reward can be higher than that under the volatile reward model. Note that this conclusion is based on the assumption of no or rare whale transactions. The presence of frequent whale transactions with substantial fees that do not follow the time-fee function pattern may offset the lower profitability of the attack in the volatile model resulting from the concavity of the time-fee function.

### v2

@@ 2.3. Non-profitability of Pre-DAM Selfish Mining Under the Fixed Block Reward Model
Under a fixed block reward model, selfish mining cannot be more profitable than honest mining during the first epoch of the attack and before the DAM is applied (Grunspan and Pérez-Marco, 2018). To compare profits from honest and selfish mining before difficulty adjustment, let $e$ denote an epoch, $\lambda$ the mining rate, $\alpha$ the mining share of adversary $\mathcal{A}$, and $R$ the fixed reward per block. We define the adversary’s total and canonical block generation rates as the average number of rounds in which a new adversarial block is mined, and the average number of rounds in which a new adversarial block is added to the canonical chain, respectively. Under honest mining, the adversary’s canonical block generation rate is $\lambda\alpha$, yielding a time-averaged profit of $\lambda\alpha R$. If the adversary $\mathcal{A}$ starts selfish mining at the beginning of epoch $e$, its canonical block generation rate would be less than or equal to that under honest mining ($\leq\lambda\alpha$). This is because the difficulty of epoch $e$ is already set before the start of the epoch, and the adversary’s strategy during the epoch cannot influence the epoch difficulty. Since both the epoch difficulty and the adversary’s mining share are fixed, the adversary’s total block generation rate remains the same under both honest and selfish mining. However, under selfish mining, some of the adversarial blocks may become orphaned, leading to a decrease in the adversary’s canonical block generation rate. Therefore, the time-averaged profit from selfish mining during epoch $e$ (before a DAM is applied) is $\leq\lambda\alpha R$. Intuitively, before a DAM is applied, the adversary cannot mine blocks faster, regardless of the strategy chosen (Sarenche et al., 2024b).
@@ 2.4. Profitability of Long-Range Selfish Mining Under the Fixed Block Reward Model
The adversary needs to sustain the selfish mining attack for a longer period to achieve profitability. In Bitcoin, DAM is responsible for adjusting the block generation rate to ensure that it takes, on average, 10 minutes to mine a new block. This implies that the ideal epoch duration is two weeks. Assume the adversary starts selfish mining at the beginning of epoch $e$. As a result of selfish mining, some of the blocks (both adversarial and honest) mined during epoch $e$ may become orphaned. The generation of these orphan blocks extends the duration of epoch $e$ because to add 2016 blocks to the canonical chain, more than 2016 blocks need to be mined. At the end of epoch $e$, which exceeds two weeks, the DAM decreases the mining difficulty for the following epoch to restore the canonical block generation rate to one block per 10 minutes. This reduction in difficulty leads to a higher block generation rate for the adversary, indicating that starting from epoch $e+1$, the adversary can mine new blocks in fewer rounds and consequently increase its time-averaged profit (Sarenche et al., 2024a). This implies that under the fixed block reward model, selfish mining requires at least one difficulty epoch to become profitable. In practice, the attack duration must be extended further to ensure profitability. This is because the selfish miner incurs losses during the initial epoch due to orphaned blocks, necessitating a longer attack period to compensate for these losses.
3. Revisiting the WeRLman Model
The WeRLman paper (Bar-Zur et al., 2023) studied selfish mining under the volatile block reward model, assuming that mining difficulty is well-adjusted during the attack. The WeRLman authors showed that the security threshold of Bitcoin mining, defined in Definition 2.5, is lower in the volatile reward model compared to the fixed block reward model. This is because, in the volatile block reward model, some blocks may occasionally be mined that include significantly higher rewards due to the inclusion of valuable transaction fees. Once such a block is mined, the adversary is incentivized to orphan it rather than extend it, in an attempt to steal the valuable transactions contained within. Although transitions to the volatile reward model can reduce the mining power threshold for profitable selfish mining, we argue that, due to the WeRLman environment design, it underestimates the mining power threshold.
The WeRLman environment: In the WeRLman environment, two types of transactions exist in the mempool: normal transactions with a normalized fee of $1$ and whale transactions with a normalized fee of $1+F$. The model assumes that, at each event of mining a block $B$ with a new height, a single transaction is added to the mempool. This transaction implicitly represents the transactions that arrive in the mempool between the mining time of block $B$ and the next block. More specifically, once a block is mined, the WeRLman environment samples a new transaction and adds it to the mempool. This new transaction is either normal, with probability $1-p$, or a whale, with probability $p$. This implementation assumption in the WeRLman model ensures a constant transaction arrival rate but introduces a significant bottleneck that benefits the adversary. In the WeRLman model, at each decision-making point, the adversary knows whether the next transaction added to the mempool will be a whale or a normal transaction before deciding its next action. We refer to this property of knowing the type of the next transaction in advance as predictive capability.
Our analysis shows that the implicit predictive capability in the WeRLman model allows the adversary to make more profitable decisions regarding whether to publish or withhold its blocks, thereby increasing its profitability compared to a real-word miner. Note that, in reality, miners do not have advance knowledge about the nature of future transactions. While mining a new block, miners frequently update the contents of their blocks to include the highest-fee transactions available. Therefore, once a block is mined, it includes the latest available whale transactions, and the miner is unaware of the arrival time of the next whale transaction. To determine Bitcoin’s security threshold in the presence of real-world adversaries, we modified the transaction sampling method in the WeRLman environment to eliminate the effect of predictive capability. We refer to this modified version of the WeRLman environment as the non-predictable WeRLman, which is explained below.
The non-predictable WeRLman environment: The only difference between the non-predictable WeRLman environment and the original version is that, in the non-predictable environment, the mempool includes only the transactions that arrived up to the block generation event, excluding future transactions. At each event of mining a block $B$ with a new height, the non-predictable environment samples a single transaction that resembles those transactions that arrived between the generation of $B$’s parent block and the generation of $B$, allowing $B$ to include the transaction. The implementation details of the non-predictable WeRLman environment are provided in Appendix C.
The authors of the WeRLman paper implemented an MDP tool to analyze the security threshold under the WeRLman environment. For a given adversary with specific mining power, this tool determines the optimal strategy for the adversary within the WeRLman environment. The security threshold is then identified as the minimum mining power at which the optimal strategy deviates from the honest strategy. This MDP tool can also be adapted to derive the optimal strategy for an adversary operating under the non-predictable WeRLman environment. In Figure 2, we compare the security threshold for a profitable deviation of honest mining in the WeRLman environment and the non-predictable WeRLman environment. To depict Figure 2, we use the same extra fee values $F$ as estimated in the WeRLman paper [^3] In the WeRLman paper, the authors analyzed Bitcoin blockchain data to estimate the extra fee value $F$ for a whale transaction, assuming whale transactions arrive with a frequency of $p=0.001$. Based on the current protocol reward of 3.125 BTC, and the protocol rewards after 1, 2, 3, 4, and 7 halvings, the estimated values of $F$ are $0.26$, $0.45$, $0.74$, $1.14$, $1.58$, and $3.2$, respectively.. Note as Bitcoin experiences more halving events, the protocol reward decreases, causing the ratio of a whale-included block reward to a normal block reward, i.e., $1+F$, to increase.
Figure 2. Bitcoin security threshold in upcoming years
As can be seen in Figure 2, the security threshold in the non-predictable environment is higher than in the WeRLman environment.
Reason for Higher Security Threshold in the non-predictable WeRLman: To explain the reason behind the higher security threshold in the non-predictable environment, we define three adversarial strategies: $\pi_{1}^{\text{WeRLman}}$, $\pi_{1}^{\text{Non-predictable}}$, and $\pi_{2}^{\text{Non-predictable}}$. The first strategy applies only in the original WeRLman environment, as it leverages predictive capability, while the other two are suited for the non-predictable WeRLman environment, where no predictive capability exists. We analyze the profitability of these strategies to identify the factors limiting the adversary’s profitability in the non-predictable setting. These three strategies are precisely defined in the Appendix D, and the main idea behind these strategist are defined as follows:
Strategy $\pi_{1}^{\text{WeRLman}}$: This strategy applies exclusively to the WeRLman environment. Once the adversary mines a normal block and knows that the next transaction added to the mempool is a whale transaction, it withholds the block and creates a private fork.  [^4] By following this strategy, the adversary creates a private fork that increases its chances of winning the whale transaction available in the mempool.. In other scenarios, the adversary follows honest behavior.
Strategy $\pi_{1}^{\text{Non-predictable}}$: This strategy is suitable for a non-predictable WeRLman environment. Once the adversary mines a normal block, it withholds the block and creates a private fork. In other scenarios, the adversary follows honest behavior.
Strategy $\pi_{2}^{\text{Non-predictable}}$: This strategy is suitable for a non-predictable WeRLman environment. Once the honest miners mine a whale-included block, the adversary does not accept the block but continues mining on the parent of the published block to undercut it. In other scenarios, the adversary follows honest behavior.
In the Appendix D, we theoretically analyze the three introduced strategies and determine their profitability. In Figure 3, we compare the security threshold under the original WeRLman and non-predictable WeRLman environments with the security thresholds of the introduced strategies: $\pi_{1}^{\text{WeRLman}}$, $\pi_{1}^{\text{Non-predictable}}$, and $\pi_{2}^{\text{Non-predictable}}$. This comparison is based on different values of $F$, a whale transaction frequency of $p=0.001$, and a security parameter $\epsilon=10^{-6}$. The security thresholds for the original WeRLman and non-predictable WeRLman environments are obtained using the MDP tool from the WeRLman paper, while the security thresholds for the introduced strategies are derived from the theoretical analysis presented in Section D. The comparison shows that the security threshold in the original WeRLman environment overlaps with the security threshold of strategy $\pi_{1}^{\text{WeRLman}}$, while the security threshold in the non-predictable WeRLman environment overlaps with the minimum of the security thresholds of $\pi_{1}^{\text{Non-predictable}}$ and $\pi_{2}^{\text{Non-predictable}}$. We can conclude that strategy $\pi_{1}^{\text{WeRLman}}$ is the deviating strategy that determines the security threshold in the original WeRLman environment. However, since strategy $\pi_{1}^{\text{WeRLman}}$ is not applicable to the non-predictable environment, the security threshold in the non-predictable environment cannot be as low as that in the original WeRLman environment.
Figure 3. Security threshold of adversarial strategies
Consider the scenario in which the adversary has mined a normal block. Under the WeRLman environment, the adversary has access to additional information about the next transaction type and can choose to withhold its normal block only in scenarios where the next transaction is known to be a whale (strategy $\pi_{1}^{\text{Non-predictable}}$). If the honest miners mine the next block and include the whale transaction, the withheld adversarial block can be used to orphan the honest block and revive the whale transaction. In other words, the adversary under the WeRLman model can withhold its normal block only if it knows that winning a fork race will guarantee the inclusion of the valuable whale transaction. However, under the non-predictable environment, if an adversary with limited mining power mines a new block, it would typically publish the block immediately to secure the reward, as it has no information about the next transaction type. This means that for adversaries with limited mining power, following strategy $\pi_{1}^{\text{Non-predictable}}$ is not profitable under the realistic non-predictable environment.